A sound wave radiates with 8.5X10^-5W of power. What is the intensity of the sound wave at a distance of .015 meters
1
Expert's answer
2020-10-29T07:01:36-0400

"I=\frac{P}{A}=\frac{P}{4\pi R^2}\\I=\frac{8.5\cdot10^{-5}}{4\pi (0.015)^2}=0.03\frac{W}{m^2}"